OpenStudy (anonymous):

What is the solution to the equation1/3(9x-6)=2x+3 ? x =

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{3}(9x-6)=2x+3 }\) like this?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

alright, can you expand the left side for me?

OpenStudy (solomonzelman):

what is: 1. \(\large\color{black}{ \displaystyle \frac{1}{3} \times 9x }\) 2. \(\large\color{black}{ \displaystyle \frac{1}{3} \times (-6) }\) ?

OpenStudy (anonymous):

I don't know what x is I have been trying but I can't

OpenStudy (solomonzelman):

I just want you to answer the 2 questions I asked you above (for now).

OpenStudy (solomonzelman):

can you simplify #1 and #2 ?

OpenStudy (anonymous):

1 is? 2 is -2

OpenStudy (solomonzelman):

2 is -2, that is correct.

OpenStudy (anonymous):

1 is3x

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \displaystyle \frac{1}{3}(9x-6)=2x+3 }\),. after expanding the left side, becomes, \(\large\color{black}{ \displaystyle 3x-2=2x+3 }\)

OpenStudy (anonymous):

ok so what is x

OpenStudy (solomonzelman):

See what happens?

OpenStudy (solomonzelman):

then, you would need to subtract 2x from both sides

OpenStudy (solomonzelman):

I have: \(\large\color{black}{ \displaystyle 3x-2=2x+3 }\) subtract 2x from both sides, \(\large\color{black}{ \displaystyle 3x-2\color{red}{-2x}=2x+3\color{red}{-2x} }\)

OpenStudy (anonymous):

so is the answer 1

OpenStudy (solomonzelman):

is this making sense, how to subtract 2x from both sides?

OpenStudy (solomonzelman):

no, it is not 1. if you plug in 1 for x, yu will see why it is not 1, because then you get 1=5.

OpenStudy (solomonzelman):

first, can you tell me what you get on each side, after subtracting 2x from both sides as I did.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3x-2\color{red}{-2x}=2x+3 \color{red}{-2x}}\) \(\large\color{black}{ \displaystyle ?=?}\) what do you get?

OpenStudy (anonymous):

yeah subtract 2x from 3x and 2x from 2x then you get 1x - 2=3

OpenStudy (anonymous):

sorry I'm lagging

OpenStudy (solomonzelman):

yes, which is same as \(\large\color{black}{ \displaystyle x-2=3}\)

OpenStudy (solomonzelman):

because \(\large\color{black}{ \displaystyle1x}\) is same as \(\large\color{black}{ \displaystyle x}\)

OpenStudy (solomonzelman):

(add 2 to both sides) I am also lagging, it is alright

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x-2=3}\) \(\large\color{black}{ \displaystyle x-2\color{royalblue}{+2}=3\color{royalblue}{+2}}\) then, \(\large\color{black}{ \displaystyle x=?}\)

OpenStudy (anonymous):

so 5 right

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

x=5:)

OpenStudy (solomonzelman):

yes, ik. So I am saying that x=5 is correct. your solution for x is 5, and that is written as x=5

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{3}(9x-6)=2x+3}\) checking: \(\large\color{black}{ \displaystyle \frac{1}{3}(9\color{green}{(5)}-6)=2\color{green}{(5)}+3}\) \(\large\color{black}{ \displaystyle \frac{1}{3}(45-6)=10+3}\) \(\large\color{black}{ \displaystyle \frac{1}{3}(39)=10+3}\) \(\large\color{black}{ \displaystyle 13=13}\)

OpenStudy (solomonzelman):

any questions?

OpenStudy (anonymous):

thanks:)

OpenStudy (solomonzelman):

yw